# LLM-Enhanced Causal Discovery in Temporal Domain from Interventional Data

### v1

@@ 5.2.2. On real-world data center datasets
Table 3 presents the results on real data. It can be observed that our method not only achieves fewer C2A False but also learns a higher number of A2A True and C2C True relations. We can also find that introducing prior domain knowledge with LLM module can effectively understand the upstream and downstream relationship of variables in the system architecture, and avoid the influence of downstream variables on upstream variables, thus controlling the C2A False continues to be zero in the subsequent optimization. Besides, the standard deviation of each metric is relatively small as well. Therefore, RealTCD outperforms other approaches comprehensively in industrial scenarios by utilizing rich information from interventional data and prior domain knowledge.
@@
Known intra+inter TECDI 85.6 $\pm$ 12.46 4.00 $\pm$ 6.85 5.80 $\pm$ 3.16 0.70 $\pm$ 1.34 2.80 $\pm$ 1.32
@@ 5.3.1. Ablation studies
In this section, we conduct ablation studies to verify the effectiveness of the proposed two modules in RealTCD, and the results are shown in Table 4.
Table 4. Ablation studies of the RealTCD modules on real-world datasets. $\uparrow$ denotes the higher the better, and $\downarrow$ denotes the lower the better. The best results in terms of metric ratios are in bold.
Method All edges C2A False $\downarrow$ A2C True $\uparrow$ A2A True $\uparrow$ C2C True $\uparrow$
@@
RealTCD-intervention 304.30 $\pm$ 18.66 0.00 $\pm$ 0.00 43.30 $\pm$ 7.33 16.50 $\pm$ 1.58 14.80 $\pm$ 1.14
[…]
RealTCD-LLM 38.80 $\pm$ 1.87 0.00 $\pm$ 0.00 0.10 $\pm$ 0.32 0.00 $\pm$ 0.00 0.40 $\pm$ 0.70
[…]
We compare the ablated version ‘RealTCD-intervention’ that removes the interventional module in score-based temporal causal discovery and uses purely observational data in the same sample size. We observe that RealTCD-intervention tends to output significantly more edges making the comparisons unfair, so we calculate the ratio of the labeled edges learned to the total number of edges learned for further evaluation. We find that removing the interventional module significantly decreases the performance, especially for the metric ratio of A2A True and C2C True, which verifies the effectiveness of the proposed score-based temporal causal discovery method from interventional data.
Similarly, we compare the ablated version ‘RealTCD-LLM’ that removes the LLM-guided meta initialization mechanism. The performance on real-world datasets significantly declines when the module is omitted, even to zero in terms of the A2A True metric, indicating that the module LLM-guided meta initialization mechanism can utilize prior information inside the abundant system textual information, and has distinct boosting effects on temporal causal discovery in real-world scenarios. This module’s integration is pivotal because it primes the causal discovery process with a well-informed starting point, thereby reducing potential biases and improving the precision of causal inference.
5.3.2. Different LLM models
We also provide a variety of attempts on different LLMs and prompts with the real-world datasets (Zhou et al., 2024; Jin et al., 2024), where the first one choosing GPT-4 and prompt with implication and Chain-of-Thoughts (CoT) (Wei et al., 2022) is the one we used in our above experiments, and all results are shown in Table 5. Overall, GPT-4 is able to achieve better and more stable results.
Table 5. Performance comparisons of different LLMs and prompt techniques in RealTCD on real-world datasets. $\uparrow$ denotes the higher the better, and $\downarrow$ denotes the lower the better.
Prompt No. implication CoT Models C2A False $\downarrow$ A2C True $\uparrow$ A2A True $\uparrow$ C2C True $\uparrow$
0 GPT-4 0 60 36 36
GPT-3.5-turbo-instructor 0 8 0 0
1 GPT-4 0 68 0 0
GPT-3.5-turbo-instructor 0 8 18 0
2 GPT-4 0 18 0 12
GPT-3.5-turbo-instructor 0 32 0 0
@@ 6.1. Motivation and strengths of using LLMs
In real-world industrial systems, operations are often accompanied by extensive textual documentation such as logs, manuals, and system descriptions. These texts provide a rich source of untapped knowledge that directly relates to system behavior and operational dependencies. LLMs excel in mining such textual data for relevant information that informs causal analysis, thereby establishing an inherent connection between textual data and system operations.
In the ”LLM-guided Meta Initialization” module, we utilize LLMs to incorporate domain knowledge and system structure information presented in textual form. This approach enables us to preliminarily extract potential causal relationships as a foundation for the discovery process. We outline the irreplaceable strengths of using LLMs over traditional deep learning approaches as follows:
  • Handling Textual Information: While traditional causal discovery methods often overlook the rich textual data available in real-world systems, LLMs excel in processing and utilizing this data. This capability significantly enhances the accuracy and relevance of causal discovery outcomes. Although conventional language models can also process text, they lack the advanced capabilities of LLMs, as detailed in the next point.
  • Integration of Domain Knowledge: Our method leverages LLMs not only to assimilate user inputs (e.g. the structure of a particular system) but also to integrate the vast domain-specific knowledge embedded within the LLMs (e.g. the operation law of a system), which may cover areas unfamiliar to the users. LLMs have access to extensive data within the same domain, often possessing more domain knowledge than the users. This integration is crucial for understanding complex systems where specific domain insights are vital for accurate causal inference.
  • LLM-guided Meta-Initialization: As demonstrated in Section 5.3.1, this innovative module significantly enhances the quality of causal discovery compared to methods that do not use textual information. Traditional temporal causal discovery methods typically initialize by assuming all relationships may have causal links, then eliminate the non-existent ones, which can lead to suboptimal local solutions and high variance in results. By utilizing meta-knowledge extracted by LLMs to narrow the scope of causal discovery initially, we can significantly improve the stability and quality of the causal discovery algorithm.
The suitability of LLMs for enhancing causal discovery is also supported by a body of literature that demonstrates their effectiveness in extracting and applying domain knowledge in complex inference tasks (Ban et al., 2023b; Chen et al., 2023; Kıcıman et al., 2023; Vashishtha et al., 2023b).
@@ 6.3. Practical Implications
Though we emphasizes the application of temporal causal discovery within AIOps, it is also helpful in various fields such as finance (Kambadur et al., 2016), healthcare (Gong et al., 2023), and social sciences (Grzymala-Busse, 2011) by uncovering the dynamic interrelations between variables over time. We discuss the broad practical implications of our method elaborately in Appendix C.
@@ 7. Conclusion
In this paper, we propose the RealTCD framework, a novel approach for temporal causal discovery in AIOps that circumvents the need for interventional targets and efficiently utilizes system textual information. Our method, with the proposed score-based temporal causal discovery and LLM-guided meta-initialization, has demonstrated its superiority over the existing baselines in both simulated and real-world datasets in terms of temporal causal discovery. The findings from RealTCD not only highlight its potential in enhancing causal analysis in complex IT systems but also open avenues for further research in AI-driven operations and maintenance in practical industrial scenarios.

### v2

@@ 4.2.2. On real-world data center datasets
Table 3 presents the results on real data. Our method achieves fewer C2A False and learns more A2A True and C2C True relations. Introducing prior domain knowledge with the LLM module effectively understands the upstream and downstream relationships in the system architecture, avoiding downstream influence on upstream variables and keeping C2A False at zero in the subsequent optimization. Additionally, the standard deviation of each metric is small. Therefore, RealTCD outperforms other approaches in industrial scenarios by utilizing rich information from interventional data and prior domain knowledge from textual information.
@@
Known intra+inter TECDI 85.60 $\pm$ 12.46 4.00 $\pm$ 6.85 5.80 $\pm$ 3.16 0.70 $\pm$ 1.34 2.80 $\pm$ 1.32
@@ 4.3.1. Ablation studies
We conduct ablation studies to verify the effectiveness of the proposed 2 modules in RealTCD as in Table 4.
‘Variant1’ is ‘RealTCD w/o intervention’, which removes the interventional module and uses only observational data of the same sample size. It tends to output more edges, making comparisons unfair, so we calculate the ratio of labeled edges to total edges learned for further evaluation. Removing the interventional module significantly decreases performance, especially for the A2A True and C2C True ratios, confirming the effectiveness of our score-based temporal causal discovery from interventional data.
‘Variant2’ is ‘RealTCD w/o LLM’, which removes the LLM-guided meta initialization. Performance drops significantly without this module, even reaching zero for the A2A True metric. This shows that the LLM-guided meta initialization effectively utilizes prior information from system textual data, greatly enhancing temporal causal discovery in real-world scenarios. Integrating this module is crucial for providing a well-informed starting point, reducing biases, and improving causal inference precision.
Table 4. Ablation studies of modules on real-world dataset. $\uparrow$ denotes the higher the better, and $\downarrow$ denotes the lower the better. The best results in terms of metric ratios are in bold.
@@
RealTCD 51.7 $\pm$ 8.0 0.0 $\pm$ 0.0 1.5 $\pm$ 1.0 14.0 $\pm$ 1.6 8.4 $\pm$ 1.9
% 0.0 $\pm$ 0.0 3.1 $\pm$ 2.4 27.3 $\pm$ 2.7 16.5 $\pm$ 3.6
Variant1 304.3$\pm$18.7 0.0 $\pm$ 0.0 43.3 $\pm$ 7.3 16.5 $\pm$ 1.6 14.8 $\pm$ 1.1
% 0.0 $\pm$ 0.0 14.2 $\pm$ 1.8 5.4 $\pm$ 0.5 4.9 $\pm$ 0.4
Variant2 38.8 $\pm$ 1.9 0.0 $\pm$ 0.0 0.1 $\pm$ 0.3 0.0 $\pm$ 0.0 0.4 $\pm$ 0.7
% 0.0 $\pm$ 0.0 0.2 $\pm$ 0.7 0.0 $\pm$ 0.0 1.0 $\pm$ 1.6
4.3.2. Different LLMs and prompts
We tested different LLMs and prompts on real-world datasets (Zhou et al., 2024; Jin et al., 2024). GPT-4, especially with prompts containing implication and CoT, achieved better and more stable results. In scenarios where human ideas are unclear, zero-shot CoT can effectively leverage LLMs’ domain knowledge, providing solutions that reveal the thinking process and enhance the interpretability of causal discovery.
Table 5. Comparisons of different LLMs and prompts. $\uparrow$ denotes the higher the better, and $\downarrow$ denotes the lower the better.
Models GPT-4 GPT-3.5-turbo-instructor
Prompt No. 0 1 2 0 1 2
Implication
CoT
C2A False $\downarrow$ 0 0 0 0 0 0
A2C True $\uparrow$ 60 68 18 8 8 32
A2A True $\uparrow$ 36 0 0 0 18 0
C2C True $\uparrow$ 36 0 12 0 0 0
@@ 5.1. Motivation and Strengths of Using LLMs
In industrial systems, operations come with extensive textual documentation like logs and manuals, providing valuable insights into system behavior. LLMs excel at mining this textual data to inform causal analysis, directly connecting text and system operations. We outline the irreplaceable strengths of using LLMs in RealTCD over traditional deep learning approaches as follows:
  • Handling Textual Information: Traditional methods often ignore the rich textual data in real-world systems, while LLMs process and utilize it to enhance causal discovery accuracy. Though conventional language models can also process text, they lack the advanced capabilities of LLMs in in-context learning (Cahyawijaya et al., 2024), which provides LLMs with superior generalization and flexibility, as well as powerful zero-shot and few-shot learning abilities (Brown et al., 2020; Li et al., 2024).
  • Integration of Domain Knowledge: LLMs assimilate user inputs (e.g. the structure of a particular system) and integrate extensive domain-specific knowledge embedded within the LLMs (e.g. the operation law of a system), covering areas unfamiliar to users. This is crucial for accurate causal inference in complex systems.
  • LLM-guided Meta-Initialization: This module significantly improves causal discovery quality by using meta-knowledge from LLMs to narrow the scope initially, unlike traditional methods that start with broad assumptions, which can lead to suboptimal local solutions and high variance in results.
The suitability of LLMs for enhancing causal discovery is also supported by a body of literature that demonstrates their effectiveness in extracting and applying domain knowledge in complex inference tasks (Ban et al., 2023b; Chen et al., 2023; Kıcıman et al., 2023; Vashishtha et al., 2023a; Jiralerspong et al., 2024).
@@ 5.3. Practical Implications
In our paper, we highlight the application of temporal causal discovery within AIOps to enhance systems’ monitoring, troubleshooting, and predictive capabilities, crucial for tasks such as anomaly detection (Yang et al., 2022b), root cause analysis (Wang et al., 2023), failure prediction, and system optimization. It is also helpful in various fields such as finance (Kambadur et al., 2016), healthcare (Gong et al., 2023; Zhang et al., 2023b; Tong et al., 2024; Naik et al., 2023), and social sciences (Grzymala-Busse, 2011) by uncovering the dynamic interrelations between variables over time. We discuss the broad practical implications elaborately in Appendix C.
6. Related Work
Causal Discovery in Temporal Domain
Interventional data greatly aids in identifying causal structures (Eberhardt, 2012; Tian and Pearl, 2013; Gao et al., 2022; Zhang et al., 2023e), but designing experiments and collecting data is challenging.  Jaber et al. (2020) used a $\Psi$-Markov property for learning causal graphs with latent variables.  (Addanki et al., 2020) proposed a $p$-collider-based algorithm to recover causal graphs with minimal intervention costs.  (Brouillard et al., 2020) used interventional data and neural architectures for causality detection. (Li et al., 2023) focused on temporal causality from observational data but required intervention labels, limiting practicality. We focus on discovering temporal causal relationships without intervention labels.
LLMs for Causal Discovery
Recent works explored LLMs for causal discovery(Jiralerspong et al., 2024; Vashishtha et al., 2023b; Bhattacharjee et al., 2023; Gui and Toubia, 2023; Kasetty et al., 2024; Liu et al., 2024; Willig et al., 2022; Takayama et al., 2024; Ban et al., 2023a; Cohrs et al., 2023) and dynamic graphs(Zhang et al., 2023d; Wang et al., 2017, 2019; Zhang et al., 2023a). Ban et al. (2023b) showed LLMs’ potential in causal inference. Chan et al. (2023) assessed ChatGPT’s ability to capture temporal and causal relations. Chen et al. (2023) used LLM-driven knowledge to reduce biases in causal learning. (Kıcıman et al., 2023) highlighted LLMs’ roles in causal discovery and reasoning. Long et al. (2023) investigated LLMs’ understanding of causal relationships in medical contexts. To the best of our knowledge, we are the first to utilize LLMs in the field of temporal causal discovery.
@@ 7. Conclusion
In this paper, we propose the RealTCD framework, a novel approach for temporal causal discovery in AIOps that bypasses the need for interventional targets and leverages system textual information. Our method, featuring score-based temporal causal discovery and LLM-guided meta-initialization, outperforms existing baselines on both simulated and real-world datasets. The results underscore RealTCD’s potential to enhance causal analysis in complex IT systems and suggest further research in AI-driven operations and maintenance.